OpenStudy (anonymous):

A star is estimated to have a mass of 1x10^36 kg. Assuming it to be a sphere of average radius 6.5x10^5 km, calculate the average density of the star in units of grams per cubic centimeter.

OpenStudy (anonymous):

density=mass/volume volume of sphere = (4/3)pi(r)^3= 1.15x10^33cm^3 mass=1x10^39 gram density=1x10^39/1.15x10^33 g/cm^3 i think this is right.

OpenStudy (anonymous):

i got the answer as you posted. it's right :) thank you

OpenStudy (anonymous):

np :)